What is thhe pH of 500mL of solution containing 0.0124 grams of Ca(OH)2?
Answer #1
How many moles do you have?
moles = grams/molar mass.
Solve for moles and plug into
M = moles/L
Note that you have 2 OH^- for each Ca(OH)2; therefore, multiply the molarity by 2 to find concn OH, then pOH = -log(OH) and convert to pH.
answered by: Shatar
Answer #2
0.0124/121.64= 1.02 x 10^-4g/mol

1.02 x 10^-4/ .5= 2.04 x 10^-4

2.04 x 10^-4*2=4.08x 10^-4

-log(4.08 x 10^-4)= 3.39
14-3.39=10.61
thank you so much!!!!!!
answered by: HM